When x^3 -4x^2 + k -5 is divided by x-2, the reminder is -1. Find the value of k.
Answer by CubeyThePenguin(3113)

f(x) = x^3 - 4k^2 + k - 5

f(2) = 2^3 - 4(2)^2 + k - 5 = -1

8 - 15 + k - 5 = -1

-11 + k = 0

k = 1